# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

### v4

@@ Abstract
We address the problem of extending the capabilities of vision foundation models such as DINO, SAM, and CLIP, to 3D tasks. Specifically, we introduce a novel method to uplift 2D image features into Gaussian Splatting representations of 3D scenes. Unlike traditional approaches that rely on minimizing a reconstruction loss, our method employs a simpler and more efficient feature aggregation technique, augmented by a graph diffusion mechanism. Graph diffusion refines 3D features, such as coarse segmentation masks, by leveraging 3D geometry and pairwise similarities induced by DINOv2. Our approach achieves performance comparable to the state of the art on multiple downstream tasks while delivering significant speed-ups. Notably, we obtain competitive segmentation results using generic DINOv2 features, despite DINOv2 not being trained on millions of annotated segmentation masks like SAM. When applied to CLIP features, our method demonstrates strong performance in open-vocabulary object localization tasks, highlighting the versatility of our approach. [^1] Project page: https://juliettemarrie.github.io/ludvig
@@ 1 Introduction
The field of image understanding has recently seen remarkable progress, driven by large pretrained models such as CLIP (Radford et al., 2021), DINO (Caron et al., 2021; Oquab et al., 2024), or SAM (Kirillov et al., 2023). A key factor behind their exceptional generalization capabilities lies in the vast size of their training datasets, often composed of millions or even billions of samples.
@@
3D scene representation has also advanced through machine learning approaches like NeRF (Mildenhall et al., 2021) and model fitting techniques such as Gaussian Splatting (Kerbl et al., 2023). These methods typically rely on a few dozen views of the scene captured from different angles. While the resulting reconstructions effectively capture both appearance and geometrical information, they are not directly applicable to semantic tasks, which has led to further developments.
@@
The complementarity of these two families of approaches has indeed recently been exploited by numerous methods that integrate geometry and semantics by uplifting image-level features extracted by large pretrained models into 3D NeRF or Gaussian Splatting representations. This has led to a surge in methods for tasks such as language-guided object retrieval (Kerr et al., 2023; Liu et al., 2023; Zuo et al., 2024), scene editing (Kobayashi et al., 2022; Chen et al., 2024; Fan et al., 2023), or semantic segmentation (Cen et al., 2023c; Ye et al., 2024a; Ying et al., 2024).
@@
The main limitation of most previous approaches lies in their dependence on optimization, which requires an iterative process to learn a scene-specific 3D representation by minimizing a reprojection error across all training views. While this loss function is intuitive, a faster and more straightforward method for transferring 2D generic visual features to already trained Gaussian splatting 3D models would be preferable, which is the purpose of this work.
@@
In this paper, we demonstrate that a simple, learning-free process is highly effective for uplifting 2D features or semantic masks into 3D Gaussian Splatting scenes. This process, which can be viewed as an ‘inverse rendering’ operation, is both computationally efficient and adaptable to any feature type. We showcase its effectiveness by uplifting visual features from DINOv2 (Oquab et al., 2024; Darcet et al., 2024), semantic masks from SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024), and language features from CLIP (Ilharco et al., 2021). Then, we show that a graph diffusion mechanism (Kondor & Lafferty, 2002; Smola & Kondor, 2003) is helpful for feature refinement in 3D scenes. This mechanism is rooted in spectral graph theory and used in spectral clustering techniques (Belkin & Niyogi, 2001; Shi & Malik, 2000; Meila & Shi, 2000). In the context of our work, it transforms coarse segmentation inputs, such as scribbles or alignment scores between visual features and a text query, into accurate 3D segmentation masks without the need for segmentation models such as SAM. When evaluated on segmentation and open-vocabulary object localization, our method achieves results comparable to state-of-the-art techniques while being significantly faster than previous approaches relying on optimization.
@@
To summarize, our contributions are threefold: (i) we introduce a simple, learning-free uplifting approach that can be directly integrated into the rendering process, achieving state-of-the-art results when applied to SAM-generated semantic masks (Sec.4.1). (ii) we demonstrate that using graph diffusion based on uplifted DINOv2 features yields competitive results for foreground/background and open-vocabulary object segmentation, despite DINOv2 not being explicitly trained for segmentation like SAM (Secs. 4.14.2). (iii) we show that combining SAM with graph diffusion achieves state-of-the-art results on open-vocabulary object segmentation tasks (Sec.4.2).
@@ 2 Related work
Learning 3D semantic scene representations with NeRF.
@@
NeRF (Mildenhall et al., 2021) uses a multilayer perceptron to predict the volume density and radiance for any given 3D position and viewing direction. Such representation can naturally be extended to semantic features. The early works N3F (Tschernezki et al., 2022) and DFF (Kobayashi et al., 2022) distill DINO 2D (i.e., image-level) features (Caron et al., 2021) in scene-specific NeRF representations. Kobayashi et al. (2022) also distill LSeg (Li et al., 2022) a CLIP-inspired language-driven model for semantic segmentation. Shortly after, LERF (Kerr et al., 2023) and 3D-OVS (Liu et al., 2023) learned 3D CLIP (Radford et al., 2021) and DINO (Caron et al., 2021) features jointly for open-vocabulary segmentation. These works were extended to other pretrained models such as latent diffusion models (Ye et al., 2023) or SAM (Kirillov et al., 2023) for semantic segmentation (Cen et al., 2023c; Ying et al., 2024).
Learning 3D semantic scene representations with Gaussian splatting.
@@
Subsequent work has relied on the more recent Gaussian splatting method (Kerbl et al., 2023), achieving high-quality novel-view synthesis while being orders of magnitude faster that NeRF-based models. Several tasks have been addressed such as semantic segmentation using SAM (Cen et al., 2023b; Ye et al., 2024a; Kim et al., 2024), language-driven retrieval or editing using CLIP combined with DINO (Zuo et al., 2024) or SAM (Ye et al., 2023), scene editing using diffusion models (Chen et al., 2024; Wang et al., 2024), and 3D-aware finetuning (Yue et al., 2024). These works learn 3D semantic representations by minimizing a reprojection loss. As a single scene can be represented by over a million Gaussians, such optimization-based techniques have strong memory and computational limitations. To handle these, FMGS (Zuo et al., 2024) employs a multi-resolution hash embedding (MHE) of the scene for uplifting DINO and CLIP representations, Feature 3DGS (Zhou et al., 2024) learns a $1\times 1$ convolutional upsampler of Gaussians’ features distilled from LSeg and SAM’s encoder, and LangSplat (Qin et al., 2024) learns an autoencoder to reduce CLIP feature dimension from 512 to 3. In contrast, our approach requires no learning, which significantly speeds up the uplifting process and reduces the memory requirements.
Direct uplifting of 2D features into 3D.
@@
Direct uplifting from 2D to 3D has been explored in prior works. OpenGaussian (Wu et al., 2024) proposes an instance-level 3D-2D feature association method that links 3D points to 2D masks, while GaussianEditor (Chen et al., 2024) uplifts 2D SAM masks into 3D to selectively optimize Gaussians for editing tasks (see details in Appendix Sec.C.3). Both approaches are specific to segmentation and are ill-suited for uplifting generic representations. Semantic Gaussians (Guo et al., 2024) pairs 2D pixels with 3D Gaussians along each pixel’s ray based on depth information but relies on a learned 3D convolutional network. In contrast, our approach is both applicable to any 2D representation and entirely parameter-free.
Leveraging 3D information to better segment in 2D.
@@
Most prior works focusing on semantic segmentation leverage 2D models specialized for this task. The early work of Yen-Chen et al. (2022) uplifts learned 2D image inpainters by optimizing view consistency over depth and appearance. Subsequent works have mostly relied on uplifting either features from SAM’s encoder (Zhou et al., 2024), binary SAM masks (Cen et al., 2023c; b), or SAM masks automatically generated for all objects in the image (Ye et al., 2024a; Ying et al., 2024; Kim et al., 2024). The latter approach is computationally expensive, as it requires querying SAM on a grid of points over the image. It also requires matching inconsistent mask predictions across views, with e.g. a temporal propagation model (Ye et al., 2024a) or a hierarchical learning approach (Kim et al., 2024), which introduces additional computational overhead. In this work, we focus on single instance segmentation and show that our uplifted features are on par with the state of the art (Cen et al., 2023c; b; Ying et al., 2024). Standing out from prior work uplifting DINO features (Tschernezki et al., 2022; Kobayashi et al., 2022; Goel et al., 2023; Kerr et al., 2023; Liu et al., 2023; Ye et al., 2023; Zuo et al., 2024), we quantitatively show that DINOv2 features can be used on their own for semantic segmentation and rival SAM-based models through a simple graph diffusion process that leverages 3D geometry.
Learning 3D CLIP features for open-vocabulary object localization.
@@
For learning 3D CLIP features, prior works also leverage vision models such as DINO or SAM. DINO is used to regularize and refine CLIP features (Kerr et al., 2023; Liu et al., 2023; Zuo et al., 2024; Shi et al., 2024), while SAM is employed for generating instance-level CLIP representations (Qin et al., 2024). These approaches suffer from high computational costs, resorting to either dimensionality reduction or efficient multi-resolution embedding representations, and usually run for a total of one to two hours for feature map generation and 3D feature optimization. In contrast, our approach bypasses the high computational cost of gradient-based optimization and, combined with graph diffusion, is an order of magnitude faster than these prior works.
@@ 3 Uplifting 2D visual representations into 3D
In this section, we present a simple yet effective method for lifting 2D visual features into 3D using Gaussian splatting and discuss its relation with more expensive optimization-based techniques.
@@ 3.1 Background on Gaussian splatting
Scene representation.
The Gaussian splatting method consists in modeling a 3D scene as a set of $n$ Gaussians densities ${\mathcal{N}}_{i}$, each defined by a mean $\mu_{i}$ in $\mathbb{R}^{3}$, a covariance $\Sigma_{i}$ in $\mathbb{R}^{3\times 3}$, an opacity $\sigma_{i}$ in $(0,1)$, and a color function $c_{i}({d})$ that depends on the viewing direction $d$ [^2] In this work, the viewing direction $d$ refers to the full camera pose, defined by its extrinsic and intrinsic parameters..
A 2D frame at a given view is an image $\hat{I}_{d}$ rendered by projecting the 3D Gaussians onto a 2D plane, parametrized by the viewing direction $d$. This projection accounts for the opacity of the Gaussians and the order in which rays associated with each pixel pass through the densities. More precisely, a pixel $p$ for a view $d$ is associated to an ordered set ${\mathcal{S}}_{{d},p}$ of Gaussians and its value is obtained by their weighted contributions:
@@
The above weights are obtained by $\alpha$-blending, i.e. $w_{i}(d,p)=\alpha_{i}(d,p)\prod_{j\in\mathcal{S}_{d,p},j<i}\left(1-\alpha_{j}(d,p)\right)$, where the Gaussian contributions $\alpha_{i}(d,p)$ are computed by multiplying the opacity $\sigma_{i}$ by the Gaussian density ${\mathcal{N}}_{i}$ projected onto the 2D plane at pixel position $p$.
Scene optimization.
Let $I_{1},\dots,I_{m}$ be a set of 2D frames from a 3D scene and $d_{1},\dots,d_{m}$ the corresponding viewing directions. Gaussian Splatting optimizes the parameters involved in the scene rendering function described in the previous section. This includes the means and covariances of the Gaussian densities, their opacities, and the color function parametrized by spherical harmonics. Denoting by $\theta$ these parameters, the following reconstruction loss is used
@@
where $\hat{I}_{d_{k},\theta}$ is the rendered frame of the scene in the direction $d_{k}$, as in Eq. (1), by using the parameters $\theta$, and $\mathcal{L}$ is a combination of $\ell_{1}$ and SSIM loss functions (Kerbl et al., 2023).
@@ 3.2 Uplifting 2D feature maps into 3D
Given a set of $m$ 2D training frames and the corresponding 3D scene obtained by Gaussian Splatting, our goal is to compute generic features $f_{i}$ in $\mathbb{R}^{c}$ for each Gaussian $i$, which would be effective for solving future downstream tasks, e.g., high-resolution semantic segmentation for new frames of the scene, or open-vocabulary object localization. In other words, $f_{i}$ can be seen as an extension of the color function $c_{i}$, even though, for simplicity, we do not consider view-dependent features in this work.
A natural approach is to consider a pretrained vision model that provides 2D feature maps for each of the $m$ frames used in Gaussian Splatting, and then devise a technique to uplift these 2D feature maps into 3D. This uplifting principle can also be directly applied to semantic masks instead of generic features, as demonstrated in Section 5. Interestingly, once the features $f_{i}$ are computed for each Gaussian $i$, it is possible to render two-dimensional feature maps for any new view, at a resolution that can be much higher than the feature maps computed for the $m$ training frames.
Uplifting with simple aggregation.
We construct uplifted features for each 3D Gaussian of the 3D Gaussian Splatting scene as a weighted average of 2D features from all frames. Each 2D feature $F_{d,p}$ from a frame at a given viewing direction $d$ and pixel $p$ contributes to the feature $f_{i}$ by a factor proportional to the rendering weight $w_{i}(d,p)$, if the Gaussian $i$ belongs to the ordered set $\mathcal{S}_{d,p}$ associated to the view/pixel pair $(d,p)$. Denoting $S_{i}=\{(d,p),i\in\mathcal{S}_{d,p}\}$ the set of view/pixel pairs contributing to the feature $f_{i}$, the resulting features are defined as follows:
$f_{i}=\sum_{(d,p)\in\mathcal{S}_{i}}\bar{w}_{i}(d,p)F_{d,p}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{w}_{i}(d,p)=\frac{w_{i}(d,p)}{\sum_{(d,p)\in\mathcal{S}_{i}}w_{i}(d,p)},$ (3)
We can interpret this equation as a normalized version of the transposed rendering operation over the $m$ viewing directions. More precisely, the rendering of any view-independent collection of features ${\bf{f}}=(f_{i})$ attached to the $n$ Gaussians into the $m$ training frames can be represented as a linear operator $W$ acting on the collection $\bf{f}$ and returning a collection of 2D feature maps ${\bf{\hat{F}}}=(\hat{F}_{d,p})$, see Eq. (4) below. Here, non-zero entries of the matrix $W$ consists of all rendering weights $w_{i}(d,p)$ when $(d,p)\in\mathcal{S}_{i}$ is placed at row $(d,p)$ and column $i$, and $\mathbf{\hat{F}}$ is a 2D matrix containing all (flattened) 2D feature maps generated for all cameras poses, with $\mathbf{\hat{F}}_{d,p}$ the feature of pixel $p$ from view at direction $d$. Similarly, the uplifting expression introduced in Eq. (3) can be expressed in terms of the transpose of $W$ and a diagonal matrix $D$ of size $m$ representing the normalization factor and whose diagonal elements are obtained by summing over the rows of $W$ as in Eq. (5) below:
@@
It is important to note that $W$ and $D$ are not explicitly constructed. Instead, they are computed by calling the forward rendering function for Gaussian Splatting and replacing the color vectors by the feature vectors. All these operations are performed within the CUDA rendering process and are illustrated in Figure 1.
Figure 1: Illustration of the inverse and forward rendering. In the inverse rendering (or uplifting) phase, features ${\bf{f}}$ are created for each 3D Gaussian by aggregating coarse 2D features ${\bf F}$ over all viewing directions. For forward rendering, the 3D features ${\bf{f}}$ are projected on any given viewing direction as in regular Gaussian Splatting. The rendering weight $\bar{w}_{i}(d,p)$ represents the relative influences of the Gaussian $i$ on pixel $p$, defined in Eq. (3).
Connection with optimization-based inverse rendering.
An alternative approach to uplifting 2D features ${\bf F}$ is to minimize a reconstruction objective $\mathcal{L}({\bf f})$, where the goal is to find uplifted features ${\bf f}$ whose rendering closely matches the original 2D features ${\bf F}$ (Tschernezki et al., 2022; Kerr et al., 2023; Zuo et al., 2024). A natural choice is to minimize the mean squared error between the 2D features ${\bf F}$ and the rendered ones ${\bf\hat{F}}$ as defined by Eq. (4):
@@
Gaussian filtering
The normalization term $\beta_{i}=\sum_{d,p\in\mathcal{S}_{i}}w_{i}(d,p)$ serves as an estimator of the relative importance of each Gaussian in the scene. Therefore, it can be used as a criterion to prune the set of Gaussians for memory efficiency. In our experiments, we filter out half of the Gaussians based on $\beta_{i}$ and observe no qualitative nor quantitative degradation of the results. This approach is inspired by prior work on efficient Gaussian Splatting representation such as proposed by Fan et al. (2023) that also prunes Gaussians based on their contribution to each pixel in the training frames.

### v5

@@ Abstract
We address the problem of extending the capabilities of vision foundation models such as DINO, SAM, and CLIP, to 3D tasks. Specifically, we introduce a novel method to uplift 2D image features into Gaussian Splatting representations of 3D scenes. Unlike traditional approaches that rely on minimizing a reconstruction loss, our method employs a simpler and more efficient feature aggregation technique, augmented by a graph diffusion mechanism. Graph diffusion refines 3D features, such as coarse segmentation masks, by leveraging 3D geometry and pairwise similarities induced by DINOv2. Our approach achieves performance comparable to the state of the art on multiple downstream tasks while delivering significant speed-ups. Notably, we obtain competitive segmentation results using only generic DINOv2 features, despite DINOv2 not being trained on millions of annotated segmentation masks like SAM. When applied to CLIP features, our method demonstrates strong performance in open-vocabulary object segmentation tasks, highlighting the versatility of our approach. [^1] Project page: https://juliettemarrie.github.io/ludvig
@@ 1 Introduction
Figure 1: In this paper, we propose a simple, parameter-free method to uplift any 2D visual features (e.g., CLIP or DINO) into a 3D Gaussian Splatting (3D-GS) representation. Uplifting is directly implemented into the rendering process, and takes about 2ms per image and feature dimension. The uplifted features can be leveraged for various 3D tasks, such as segmentation or object removal, based on various inputs, such as scribbles or text prompts.
@@
Image understanding has made remarkable progress, driven by large pretrained models such as CLIP [36], DINO [2, 33], or SAM [21], also called foundation models. A key factor behind their exceptional generalization capabilities lies in the size of their training datasets, often composed of millions or even billions of samples.
@@
Meanwhile, 3D scene representation has also advanced through machine learning approaches like NeRF [31] and model fitting techniques such as Gaussian Splatting [18]. These methods typically reconstruct scenes from a few dozen views captured from different angles, effectively preserving both appearance and geometric details. However, they are not suited for semantic tasks.
@@
The complementarity of these two families of approaches has been exploited to integrate geometry and image-level semantics extracted by large 2D pretrained models into NeRF or Gaussian Splatting 3D representations. This has led to a surge in methods for language-guided object retrieval [19, 28, 56, 46], scene editing [23, 6, 10], or semantic segmentation [4, 50, 52] in 3D scenes.
@@
The main limitation of these approaches lies in their dependence on optimization, which requires an iterative process to learn a scene-specific 3D representation by minimizing a reprojection error across all training views. While this loss function is intuitive, a faster and more straightforward method for transferring 2D generic visual features to already trained Gaussian Splatting 3D models would be preferable, which is one of the purposes of this work.
@@
In this paper, we demonstrate that a simple, learning-free process is highly effective for uplifting 2D features or semantic masks into 3D Gaussian Splatting scenes. This process, which can be viewed as an ‘inverse rendering’ operation, is both computationally efficient and adaptable to any feature type. We showcase its effectiveness by uplifting visual features from DINOv2 [33, 9], semantic masks from SAM [21] and SAM2 [37], and visual features from CLIP [15], enabling language-driven applications. Then, we show that a graph diffusion mechanism [24, 42] is helpful for feature refinement in 3D scenes. This mechanism is rooted in spectral graph theory and spectral clustering [1, 40, 29]. In the context of our work, it transforms coarse segmentation inputs, such as scribbles or alignment scores between visual features and a text query, into accurate 3D segmentation masks without the need for segmentation models such as SAM. When evaluated on segmentation and open-vocabulary object localization tasks, our method matches state-of-the-art performance while being significantly faster than previous optimization-based approaches.
@@
To summarize, our contributions are threefold: (i) We introduce a simple, learning-free uplifting approach that can be directly integrated into the rendering process (Sec. 3.2), achieving state-of-the-art segmentation results when applied to SAM-generated semantic masks (Sec. 4.1); (ii) We introduce a graph diffusion process for 3D segmentation based on uplifted features (Sec. 3.3) which yields competitive results in both foreground/background segmentation with DINOv2 alone and open-vocabulary segmentation when combined with CLIP (Secs. 4.1, 4.2); (iii) We show that our simple uplifting approach can replace more complex and time-consuming feature learning methods and yield strong accuracy gains (Sec. 4.3).
@@ 2 Related work
Learning 3D semantic scene representations with NeRF. NeRF [31] uses a multilayer perceptron to predict the volume density and radiance for any given 3D position and viewing direction. Such a representation can naturally be extended to semantic features. The early works N3F [43] and DFF [23] distill DINO 2D (i.e., image-level) features [2] in scene-specific NeRF representations. Kobayashi et al. [23] also distill LSeg [26], a language-driven model for semantic segmentation. Building on this, LERF [19] and 3D-OVS [28] learn 3D CLIP [36] and DINO [2] features jointly for open-vocabulary segmentation. These works were extended to other pretrained models such as latent diffusion models [49] or SAM [21] for segmentation [4, 52].
@@
Learning 3D semantic scene representations with GS. Subsequent work has relied on the Gaussian Splatting method [18], enabling rendering speeds orders of magnitude faster than NeRF-based models. Several tasks have been addressed such as semantic segmentation using SAM [3, 50, 20, 46], language-driven retrieval or editing using CLIP combined with DINO [56] or SAM [49, 46], scene editing using diffusion models [6, 44], and 3D-aware finetuning [53]. These works learn 3D semantic representations by minimizing a reprojection loss. As a single scene can be represented by over a million Gaussians, such optimization-based techniques have strong memory and computational limitations. To handle these, FMGS [56] employs a multi-resolution hash embedding (MHE) of the scene for uplifting DINO and CLIP representations, Feature 3DGS [55] learns a $\times 1$ 1 convolutional upsampler of Gaussians’ features distilled from LSeg and SAM’s encoder, and LangSplat [35] learns an autoencoder to reduce CLIP feature dimension from 512 to 3. In contrast, our approach requires no learning, which significantly speeds up the uplifting process and reduces the memory requirements.
@@
Direct uplifting of 2D features into 3D. Direct uplifting from 2D to 3D has been explored in prior works. GaussianEditor [6] uplifts 2D SAM masks into 3D to selectively optimize Gaussians for editing tasks, see details in the supplementary (Sec. D.4). Dr. Splat [17] also lifts 2D SAM masks, to create an assignment from object CLIP features to Gaussians. Both approaches are specific to segmentation and are ill-suited for uplifting generic representations. Semantic Gaussians [13] pairs 2D pixels with 3D Gaussians along each pixel’s ray based on depth information but relies on a learned 3D convolutional network. OpenGaussian [46] also pairs pixels with Gaussians, but relies on pseudo-features learned with a contrastive loss on SAM masks. In contrast, our approach is simple to implement, applicable to any 2D representation and entirely parameter-free.
@@
Leveraging 3D information to better segment in 2D. Most prior works focusing on semantic segmentation leverage 2D models specialized for this task. The early work of [51] uplifts learned 2D image inpainters by optimizing view consistency over depth and appearance. Subsequent works have mostly relied on uplifting either features from SAM’s encoder [55], binary SAM masks [4, 3], or SAM masks automatically generated for all objects in the image [50, 52, 20]. The latter approach is computationally expensive, as it requires querying SAM on a grid of points over the image. It also requires matching inconsistent mask predictions across views, with e.g. a temporal propagation model [50] or a hierarchical learning approach [20], which introduces additional computational overhead. In this work, we focus on single instance segmentation and show that our uplifted features are on par with the state of the art [4, 3, 52]. Standing out from prior work uplifting DINO features [43, 23, 12, 19, 28, 49, 56], we show that DINOv2 features can be used on their own for semantic segmentation and rival SAM-based models through a simple graph diffusion process that leverages 3D geometry.
@@
3D CLIP features for open-vocabulary localization. For learning 3D CLIP features, prior works also leverage vision models such as DINO or SAM. DINO is used to regularize and refine CLIP features [19, 28, 56, 41], while SAM is employed for generating instance-level CLIP representations [35, 46, 17]. These approaches suffer from high computational costs, resorting to dimensionality reduction or efficient multi-resolution embedding representations, and usually run for a total of one to two hours for feature map generation and 3D feature optimization. In contrast, our approach bypasses the high computational cost of gradient-based optimization and, combined with graph diffusion, is an order of magnitude faster than these prior works.
@@ 3 Uplifting 2D visual representations into 3D
We now present a simple yet effective method for lifting 2D visual features into 3D using Gaussian Splatting, and discuss its relation with optimization-based techniques.
@@ 3.1 Background on Gaussian Splatting
Scene representation. The Gaussian Splatting method consists in modeling a 3D scene as a set of $n$ Gaussians densities ${\mathcal{N}}_{i}$, each defined by a mean $\mu_{i}$ in $\mathbb{R}^{3}$, a covariance $\Sigma_{i}$ in $\mathbb{R}^{3\times 3}$, an opacity $\sigma_{i}$ in (0,1)(0,1)( 0 , 1 ), and a color function $c_{i}({d})$ that depends on the viewing direction $d$. $d$ refers to the full camera pose, i.e. its extrinsics and intrinsics.
A 2D frame at a given view is an image $\hat{I}_{d}$ rendered by projecting the 3D Gaussians onto a 2D plane, parametrized by the viewing direction $d$. This projection accounts for the opacity of the Gaussians and the order in which rays associated with each pixel pass through the densities. More precisely, a pixel $p$ for a view $d$ is associated to an ordered set ${\mathcal{S}}_{{d},p}$ of Gaussians and its value is obtained by summing their contributions:
@@
The above weights are obtained by $\alpha$-blending, i.e. $w_{i}(d,p)=\alpha_{i}(d,p)\prod_{j\in\mathcal{S}_{d,p},j<i}\left(1-\alpha_{j}(d,p)\right)$, where the Gaussian contributions $\alpha_{i}(d,p)$ are given by the product of the opacity $\sigma_{i}$ and the Gaussian density $\mathcal{N}_{i}$ projected onto the 2D plane at pixel position $p$.
Scene optimization. Let $I_{1},\dots,I_{m}$ be a set of 2D frames from a 3D scene and $d_{1},\dots,d_{m}$ the corresponding viewing directions. Gaussian Splatting optimizes the parameters involved in the scene rendering function described in the previous section. This includes the means and covariances of the Gaussian densities, their opacities, and the color function parametrized by spherical harmonics. Denoting these parameters by $\theta$, the following reconstruction loss is used
@@
where $\hat{I}_{d_{k},\theta}$ is the frame of the scene rendered in the direction $d_{k}$ as in Eq. (1), using the parameters $\theta$, and $\mathcal{L}$ is a combination of $\ell_{1}$ and SSIM loss functions [18].
Figure 2: Illustration of uplifting and rendering. In the uplifting phase, features ${\bf{f}}$ are created for each 3D Gaussian by aggregating coarse 2D features ${\bf F}$ over all viewing directions. For rendering, the 3D features ${\bf{f}}$ are projected on any given viewing direction as in regular Gaussian Splatting. The rendering weight $\bar{w}_{i}(d,p)$ represents the relative influence of the Gaussian $i$ on pixel $p$ (Eq. (3)).
@@ 3.2 Uplifting 2D feature maps into 3D
Given a 3D Gaussian Splatting scene representation, we propose a method to uplift any set of 2D features maps $\bf{F}=\{F_{1},...,F_{m}\}$ associated to $m$ views into 3D. These features may be obtained from pretrained models [36, 33] or segmentation masks. The uplifted features ${\bf f}$ should allow for downstream 3D tasks, such as 3D segmentation, while being ‘faithful’ when reprojected back in 2D.
[…]
$f_{i}=\!\!\sum_{(d,p)\in\mathcal{S}_{i}}\!\bar{w}_{i}(d,p)F_{d,p},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{}\leavevmode\nobreak\ \leavevmode\nobreak\ \bar{w}_{i}(d,p)=\frac{w_{i}(d,p)}{\sum_{(d,p)\in\mathcal{S}_{i}}\!w_{i}(d,p)}.$ (3)
[…]
@@
Note that $W$ and $D$ are never explicitly constructed. Instead, they are computed by calling the forward rendering function for Gaussian Splatting and replacing the color vectors by the feature vectors (see Fig. 2). All these operations are performed within the CUDA rendering process.
Connection with optimization-based inverse rendering. An alternative approach to uplifting 2D features ${\bf F}$ is to minimize a reconstruction objective $\mathcal{L}({\bf f})$, where the goal is to find uplifted features ${\bf f}$ whose rendering closely matches the original 2D features ${\bf F}$ [43, 19, 56]. A natural choice is to minimize the mean squared error between the 2D features ${\bf F}$ and the rendered ones $^{\bf\hat{F}}$ as defined by Eq. (4):
@@
Gaussian filtering. The normalization term $\beta_{i}=\sum_{d,p\in\mathcal{S}_{i}}w_{i}(d,p)$ serves as an estimator of the relative importance of each Gaussian in the scene. Therefore, it can be used as a criterion to prune the set of Gaussians for memory efficiency. In our experiments, we filter out half of the Gaussians based on $\beta_{i}$ and observe no qualitative nor quantitative degradation of the results. This approach is inspired by prior work on efficient Gaussian Splatting representation such as proposed by [10] that also prunes Gaussians based on their contribution to each pixel in the training frames.